OpenStudy (anonymous):

km*mn*nk. I don't understand it. Help! I need somebody, Help! Not just anybody!....A maths tutor.

OpenStudy (anonymous):

No nick not you, you didn't explian properly before. Like I dont get how to do them.

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

do you need to simplify this: km*mn*nk

OpenStudy (anonymous):

peform it then simplify the answer

OpenStudy (anonymous):

when we multiply something by itself it's called "squaring" and you write the number with a small "2" in the "superscript" (above) : \[k \times m \times m \times n \times n \times k = k \times m^2 \times n^2 \times k\] so now we have \[k \times m^2 \times n^2 \times k \] ok now when you multiply things you can swap them round, the technical name for this property is "commutative" . for example you'd agree that 2x3 = 3x2 this works in algebra as well \[k \times m^2 \times n^2 \times k = k \times k \times m^2 \times n^2\] and simplifying once more: \[ k \times k \times m^2 \times n^2 = k^2 m^2 n^2\] in fact we can write the whole thing like this: \[(kmn)^2\]

OpenStudy (anonymous):

is that what you were after?

OpenStudy (anonymous):

Yes yes thankyou very much my dear friend!!!

OpenStudy (anonymous):

so you understand fine? if anything is still confusing just say and ill try to help :)

OpenStudy (anonymous):

Aw ok. I'll add you and rate you and stuff. Good maths teacher!